OpenStudy (anonymous):

Help needed. -6= -9 -3x x=

OpenStudy (anonymous):

a=8

undeadknight26 (undeadknight26):

-5 + 4a = 27 5 - 4a = 27 +5 +5 4a = 32 /4 /4 a = 8

OpenStudy (anonymous):

-5+4a=27 4a=27+5 4a=32 a=32/4 a=8

OpenStudy (anonymous):

Whats? -3+5x=42 x=

undeadknight26 (undeadknight26):

well do the same thing we did and you will get ur answer XD

OpenStudy (anonymous):